During the year, the Senbet Discount Tire Company had gross sales of $1.12 million. The company's cost of goods sold and selling expenses were $531,000 and $221,000, respectively. The company also had debt of $860,000, which carried an interest rate of 6 percent. Depreciation was $136,000. The tax rate was 40 percent.

a.
What was the company's net income? (Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars, e.g., 1,234,567. Round your answer to the nearest whole number, e.g., 32.)

     

b.
What was the company’s operating cash flow? (Do not round intermediate calculations. Enter your answer in dollars, not millions of dollars, e.g., 1,234,567. Round your answer to the nearest whole number, e.g., 32.)
Answer #1

a)Net income before tax=sales-cost of goods sold-selling expenses-depreciaiton-interest expenses-tax
=((1.12*10^6)-(531000)-(221000)-(136000)-(860000*6%))
=180400
Tax=net income before tax*tax rate
=180400*40%=72160
Net income=180400-72160=108240

b)Operating cash flow= netincome+depreciation-tax rate
=108240+136000-72160
=172080